Show that Gauss’ Law can be used to have an equation that shows the capacitance of two parallel plates in a vacuum.


1
Expert's answer
2021-02-06T10:04:24-0500

"\int EdA=\frac{1}{\epsilon_0}\sum{q_i\to E=\frac{q}{\epsilon_0A}}"


"V=Ed"


"C=\frac{q}{V}=\frac{q}{Ed}=\frac{\epsilon_0Aq}{qd}=\frac{\epsilon_0A}{d}" . Answer